Wizco Advertising's balance sheet data at May 31, 2018, and June 30, 2018, follow: 


May 31, 2018June 30, 2018
Total Assets$122,000$287,000
Total Liabilities66,000144,000


For each of the following situations that occurred in June 2018 with regard to the owner's contributions and withdrawals, compute the amount of net income or net loss during June 2018. 

a. The owner contributed $10,000 to the business and made no withdrawals. 

b. The owner made no contributions. The owner withdrew cash of $3,000. 

c. owner made contributions of $12,500 and withdrew cash of $30,000 


5. As the manager of a Papa Sean's restaurant, you must deal with a variety of business transactions. Give an example of a transaction that has each of the following effects on the accounting equation: 

a. Increase one asset and decrease another asset. 

b. Decrease an asset and decrease equity. 

c. Decrease an asset and decrease a liability. 

d. Increase an asset and increase equity. 

e. Increase an asset and increase a liability.
Answer #1

Solutions :

Assets = Equity +Liabilities
where Asset = 287500
Liabilities   = 144000
287500 = equity +140000
Equity = 147500
The Equity is woth 147500
Closing equity balance =opening Eq balance+owner contribution-owners withdrawal+ net income/loss
147500
=122000+10000-0 + net incom3/loss
Net Income 13500

Case 3

The Equity is woth 147500
Closing eq balance = opening Eq balance+owner contribution-owners withdrawal+ net income/loss
147500 = 122000+12500 -30000 net incom3/loss
Net Income /loss 43000

Case 2..

The Equity is woth 147500
Closing eq balance = opening Eq balance+owner contribution-owners withdrawal+ net income/loss
147500 = 122000-3000 + net incom3/loss
Net Income /loss 28500

Part b

A)Increase in one asset, decrease in another asset.- Purchase of Furniture  in cash ie. increase in asset and decrease in cash
b)Decrease in asset, decrease in Equity - drawings (withdrawn by Propreitor) ie, decrease in capital and decrease in cash

c)Decrease in asset, decrease in liability -Payment to creditors ie, decrease in cash and decrease in creditors